OpenStudy (ellalaynemaguire):

FAN AND MEDAL! (28 ⋅ 3−5 ⋅ 60)−2 ⋅ 3 to the power of negative 2 over 2 to the power of 3, whole to the power of 4 ⋅ 228

OpenStudy (redheadangel):

Have you tried using PEMDAS?

OpenStudy (redheadangel):

by the way, 6 ^0=1, so keep that in mind

OpenStudy (samantha2234):

mathway works

OpenStudy (redheadangel):

(2^8 × 5^–5 × 19^0)–2 ×× 228 = (2^–16 × 5^10 × 19^0) × × 2^28 = (2^–16 × 5^10 × 1) × × 2^28 = (2^–16 × 5^10 × 1) × 5^–8 × 2^–12 × 2^28 = (2^–16 × 2^–12 × 2^28) ×(5^10 ×5^–8) = 2^–28 × 2^28× 5^2 = 2^0 × 25 = 25